I THINK slope-intercept form is supposed to be used to solve this question:
At the surface of the Atlantic Ocean, the water pressure is the same as the air pressure abov the water, 101.3 KPa (Kilo Pascal). Below the surface, the water pressure increases by 29.9 KPa for every 10 ft of descent. At what depth (approximated to the first digit) is the pressure 690 KPa?
Please help. :]

Answer by mathdoc314(58)


Let's let the independent variable be d = depth in feet and the pressure is P(d) in KPa.
The information given is P(0) = 101.3 and (seems like) P is a linear function and the slope is 29.9/10.
So it looks like P(d) = 101.3 + 2.99 d
There is no other info in the problem to correct for the curvature of the earth, the decreasing distance to the center of the earth, the changing density of the water due to temperature, perturbations due to shark fins etc. so the statement of linearity is a reasonable assumption to make given the available information.
All you gotta do is solve 101.3 + 2.99d = 690
2.99d = 690 - 101.3 = 688.7
d = 688.7 / 2.99 = 230.34448160535 (according to d ~= 230.3 to 1 significant decimal place.